L^{A}TEX2ε—A New Version of L^{A}TEX

Chris Rowley and Frank Mittelbach

L^{A}TEX3 Project

This is a copy of an article in TEX and TUG news.

What is it? L^{A}TEX2ε is a new, standard version of L^{A}TEX—prepared and supported by the L^{A}TEX3 Project Team.

Why is it needed? Over the years many extensions of L^{A}TEX have been developed. This is, of course, a welcome development, since it shows that the L^{A}TEX system is in a healthy state. It has, however, had one unfortunate consequence: there are now several incompatible systems, in the sense of format (.fmt) files, all claiming to be L^{A}TEX. Therefore, in order to process documents coming from various places, a site maintainer needs to provide several format files: L^{A}TEX (with and without NFSS), SliTEX, AMS-L^{A}TEX, and so on. In addition, when looking at a source file it is not always clear for which format the document was written.

What will it offer? L^{A}TEX2ε will put an end to this unsatisfactory situation— it will give access to all such extensions based on a single format and thus end the proliferation of mutually incompatible dialects of L^{A}TEX 2.09.

It will use an enhanced version (NFSS2) of the New Font Selection Scheme. Files such as amstex.sty (formerly the AMS-L^{A}TEX format) or slides.sty (formerly the SliTEX format) will become extension packages, all working with this single format.

The introduction of this new version will also make it possible to add a small number of often-requested features (such as extended versions of \newcommand).

To summarize:

• Standardisation: a single format incorporating NFSS2, to replace the present multiplicity of incompatible formats (NFSS, lfonts, pslfonts, etc.)

• Maintenance: a standardised system supported by a reliable maintenance policy.

Compatibilty L^{A}TEX2ε will, as far as possible, adhere to the following principles:

1. Unmodified version 2.09 document files will produce the same output with L^{A}TEX2ε as with version 2.09.

2. All new features of L^{A}TEX2ε will conform to the conventions of version 2.09, making it as easy as possible for current users to learn to use them.

Documentation The new version will be described in a new edition of Leslie Lamport’s L^{A}TEX: A Document Preparation System, and in a new book, The L^{A}TEX Companion by Michael Goossens, Frank Mittelbach and Alexander Samarin (both to be published by Addison-Wesley). The Companion will also contain a complete description of NFSS2.

Distribution Policy Maintenance of the new system will be undertaken by the L^{A}TEX3 Project Team.

A complete distribution of all files, incorporating corrections of errors, will be made available twice a year on fixed dates. This will happen even if there have been no changes to the files, and hence only the release dates have had to be updated. It will therefore be easy to determine whether you have all the latest bug fixes—information which is essential to system maintainers.

Error Reports Error reports can be made using a report generating program latexbug.tex. This will be part of the main distribution.

Error reports will be accepted only if the version of L^{A}TEX2ε that produced the error is not older than one year. Error reports can be sent to this e-mail

address: latex-bugs@rus.uni-stuttgart.de.

Availability The α-test version of L^{A}TEX2ε is scheduled for late fall and the first regular distribution will be in early 1994 (says Frank). This will also contain information and advice to help in the process of updating an installation for use with the new version (says Chris).